OpenStudy (anonymous):

Simplify the expression. sine of x to the second power minus one divided by cosine of negative x

OpenStudy (anonymous):

@campbell_st @ParthKohli @IMStuck @SithsAndGiggles

OpenStudy (imstuck):

I can try...let me take a look!

OpenStudy (anonymous):

Thanks! This is the last one I have to do today :D

OpenStudy (imstuck):

Is this your expression?\[\sin ^{2}x-\frac{ 1 }{ \cos ^{-x} }\]?

OpenStudy (imstuck):

Got it. Let me study that for a tiny minute.

OpenStudy (imstuck):

Ok, looking at the numerator, we have an identity from which we can derive the simplification. Do you know your Pythgorean identities in trig?

OpenStudy (anonymous):

sin^2 θ + cos^2 θ = 1 tan^2 θ + 1 = sec^2 θ cot^2 θ + 1 = csc^2 θ

OpenStudy (anonymous):

\[ \cos(-x) = \cos(x) \]

OpenStudy (imstuck):

Ok use the first one. Get it into that same format as your numerator. Do you know what I mean? Looking at your identity, what does sin^2(x) - 1 equal?

OpenStudy (anonymous):

\[ \sin^2(x) = 1-\cos^2(x) \]

OpenStudy (imstuck):

Please let her figure it out for herself. She needs to learn.

OpenStudy (anonymous):

-cos?

OpenStudy (imstuck):

-cos^2 (x). \[\sin ^{2}x-1=-\cos ^{2}x\]That was good. Now let's replace that numerator with its identity.

OpenStudy (imstuck):

\[-\frac{ \cos ^{2}x }{ \cos(-x) }\]As wio stated up above, cos(-x) = cos(x). So we can make that replacement as well.

OpenStudy (imstuck):

\[-\frac{ \cos ^{2}x }{ cosx }\]How do you reduce/simplify that now?

OpenStudy (anonymous):

-cos/1?

OpenStudy (imstuck):

The denominator will completely cancel. Left up on top, when you factor one cos x out what are you left with?

OpenStudy (anonymous):

cos or -cos?

OpenStudy (imstuck):

Oh, I see what you did. You're right! Sorry! The 1 in the denominator threw me for a sec, as I always leave it out when it's a 1. But you're right!

OpenStudy (imstuck):

It's -cos(x)

OpenStudy (anonymous):

so that's it, right? just factor out cos and I'm left with -cosx?

OpenStudy (imstuck):

Yep!! Trig uses identities ALL the TIME! So get used to looking for a way to use one in most problems.

OpenStudy (anonymous):

Thanks again! :D

OpenStudy (imstuck):

You're most welcome!